A 500 g bar of ice is initially at -25 °C. Calculate the amount of energy required to vapourize 25% of the substance at 100 °C.


1
Expert's answer
2021-06-09T08:05:15-0400

The specific heat capacities are following:

ice: 2.093 J/goC

water: 4.184 J/goC

The heat of fusion of H2O (Lice) = 333.55 J/g

The heat of vaporization of H2O (Lsteam) = 2256 J/g


Heat required to heat the ice from -25oC to the melting (0oC):

"Q_1=c_{ice}m\Delta{T}=2.093J/g\cdot{\degree{C}}\times500g\times(0\degree{C}-(-25\degree{C}))=26163J"


Heat required to melt the ice at 0oC:

"Q_2=L_{ice}m=333.55J/g\times500g=166775J"


Heat required to heat the water from 0oC to the boiling (100oC):

"Q_3=c_{water}m\Delta{T}=4.184J/g\cdot{\degree{C}}\times500g\times(100\degree{C}-0\degree{C})=209200J"


Heat required to evaporate 25% of water (which is "500g\times0.25=125g" ) at 100oC:

"Q_4=L_{steam}m=2256J/g\times125g=282000J"


The total heat is the sum of the heats required for each step:

"Q=Q_1+Q_2+Q_3+Q_4=26163J+166775J+209200J+282000J=684138J=684kJ"


Answer: 684 kJ